It is easier to rotate a rod about its center or about its edge and why? Explain with appropriate equations
Answer #1

It is easier to rotate the rod about its center than about edge.

let M is the mass and L is the length of the rod.

moment of inertia of the rod about an axis passing through center of mass perpendicular to its length,

Icm = M*L^2/3

moment of inertia of the rod about its edge,

I_edge = M*L^2/3

clearly, I_edge = 4*Icm

so, it is easier to rotate the rod about its center.